OpenStudy (anonymous):

Solve the system of linear equations below. 4x - 3y = 17 5x + 4y = 60

OpenStudy (anonymous):

4y=60-5x => y=15-5x/4 4x-3y=17 4x-3(15-5x/4)=17 4x-45+15x/4=17 x(4+15/4)=62 x(31/4)=62 x=8 y=15-5x/4=15-5(8)/4=15-10=5